0:30 Steve restarted processes and powered up the WFS. No issues. Put the FLAO GUI in Auto for OSCO.

1:30 Steve sent pointing and collimation

1:53 Seinding OSCO HD61059. Optimize gain failed. The seeing is pretty bad so thtis may not be possible.

Still thinks the loops is closed after the preset was cancelled from AOS. Opened loop from WFS control gui which opened on Adsec and WFS. All are synced up again.

Resending preset to OSCO target. Saturating position X (y on gopt). Rotator at 363. Completed successfuly but barely. Changing PA by 50deg 303deg. Seeing conditions are not suitable for applying optical gain. Seeing blew up to 3" while we were optimizing. We gave up on OSCO since it is now 12deg and this would be useless.

-- JenniferPower - 06 Mar 2021
Topic revision: r1 - 06 Mar 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback